TipTop Limited declared that the pretax income for their company for 2019 was $6,870M. The Sales revenue and the COGS for the company were $25,020M and $10,050M respectively.

a. What is the new pretax income if the purchasing/ procurement reduces COGS by 2%?

b. How much would Sales/Marketing need to increase Sales to have the same effect as reducing COGS by 2%?
Answer #1

Answer:

  • Pretax income = A= $ 6,870 M
  • Sales = B = $25,020 M
  • COGS = C =$ 10,050 M
  • Pretx income as % of sales = D = A/B = 6870/25020 = 0.2748 =27.48%

a)

  • When there is 2% reduction in COGS, all 2% savings will increase the pre tax income directly. So :
    • the increase in pretax income = reduction in COGS
    • 2% decrease in COGS = 2% of $10050 M = $201M
    • So the increase in pretax income = E = $201 M
  • New Pretax income = A+ E = $7,071 M (Answer)

B)

  • Now pre tax income is 27.48 % of sales, so to increase pre tax income by $201 M , sales will have to increase in same ratio ( 1/0.2748). Hence:
    • Increase in sales to have $201 M increase in pre tax inxome = E /D = $ 732.0262 M (Ans)